# Path planning with moving obstacles using stochastic optimal control

### v1

@@ Abstract
Navigating a collision-free, optimal path for a robot poses a perpetual challenge, particularly in the presence of moving objects such as humans. In this study, we formulate the problem of finding an optimal path as a stochastic optimal control problem. However, obtaining a solution to this problem is nontrivial. Therefore, we consider a simplified problem, which is more tractable. For this simplified formulation, we are able to solve the corresponding Bellman equation. However, the solution obtained from the simplified problem does not sufficiently address the original problem of interest. To address the full problem, we propose a numerical procedure where we solve an optimization problem at each sampling instant. The solution to the simplified problem is integrated into the online formulation as a final-state penalty. We illustrate the efficiency of the proposed method using a numerical example.
@@ I INTRODUCTION
There are numerous applications of autonomous mobile robots that are discussed in the literature. In general, these applications are divided into two groups: indoor applications, such as delivering packages, cargo, and cleaning large buildings [1], and outdoor field robotics application [2]. Finding an optimal path for the robot to reach its destination is a crucial task in these applications. Path planning is one of the primary challenges that must be solved before mobile robots can autonomously navigate and explore complex environments [3]. Based on the nature of the environment that the robot is working in, the mobile path planning is divided into two classes, i.e. static environment and dynamic environment [1][4]. In the static environment, the obstacles are static and do not move. However, in the dynamic environment, the obstacles move during the planning task, while the robot is moving to reach its destination. The primary goal of path planning is to ensure safe, efficient, and collision-free navigation in both static and dynamic environments [5]. Path planning algorithms can be broadly categorized into several groups. The first group is classical algorithms, such as Dijkstra’s algorithm. The second group is heuristic-based approaches, e.g., the $A^{\star}$ algorithm. The third group is artificial intelligence-based approaches, such as reinforcement learning, fuzzy logic, and so on. The fourth group is hybrid-based approaches, which combine elements from all the above classes of path planning [1]. However, this is not the only way to classify path planning algorithms. In [4][6], and [7], other classifications for different approaches to path planning are also presented.
@@
In this paper, we consider a case in which there is a robot and a stochastically moving obstacle in the environment. Also, we assume that there is a static target, and we want to control the robot to reach this target. We formulate the problem of finding an optimal path as an infinite horizon stochastic optimal control problem. Numerous deterministic optimal control problem formulations are available for path planning [8]; however, to the best of the author’s knowledge, the stochastic setting has not been considered. The associated stochastic Bellman equation can be solved using value iteration. Notably, for many problems, achieving an exact solution is simply unfeasible due to what Bellman properly termed the ‘curse of dimensionality,’ in his book [9, p. 94]. This problem refers to the exponential increase in the required computations as the problem size grows [10, p. 282]. In order to circumvent this problem, we proposed a method based on geometrical symmetry to reduce the dimension of the domain of the value function. Then we use fitted value iterations to find the value function. Fitted value iterations is a parametric approximation method in which we parameterize the value function as a linear regression such that it closely satisfies the Bellman equation [11, p. 211]. Fitted value iteration is a special case of approximate value iteration, dating back to 1959 [12]. We do not explicitly compute the policy, since it is defined in the original domain of the value function. Instead, we compute the control signal in real time for the given current values of the states.
We will compare our method with a receding horizon $A^{\star}$ algorithm. $A^{\star}$ is one of the most popular path finding algorithms [13]. In this paper, we consider a receding horizon version of this algorithm. We assume that the environment is static and find the path to reach the target using the $A^{\star}$ algorithm. At each stage, we take one step and discard the rest of the path generated by the $A^{\star}$ algorithm. One of the main drawbacks of this method is its large execution time [13].
@@
    We formulate an optimal path planning problem as a stochastic optimal control problem.
@@
    We show that the corresponding value function exhibits symmetry, enabling its representation with a reduced number of variables.
  3. 3.
    We use a numerical example to illustrate that the proposed framework can outperform the widely used $A^{\star}$ algorithm on the considered class of problems.
@@ II Mathematical Model
Consider the problem where there is a robot, a stochastically moving obstacle, and a static target in a $2$-dimensional space. The goal for the robot is to reach the target without colliding with the obstacle. It should at the same time reach the target as fast as possible. These are conflicting objectives that need to be traded off. In this section, we define the mathematical model of the dynamics and formulate an optimal control problem that will meet the objectives.
@@ II-A System Dynamics
Let $r_{k}=\begin{bmatrix}r_{k}^{x}&r_{k}^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ and $h_{k}=\begin{bmatrix}h_{k}^{x}&h_{k}^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ represent the positions of the robot and the moving obstacle at time $k$, respectively. Moreover, assume that $t\in\mathbb{R}^{2}$ is the target position, where $\mathbb{R}$ is the set of real numbers. The dynamics of the robot and the moving obstacle are described by the following equations:
@@
where $F_{r}:\mathbb{R}^{2}\times\mathcal{U}\rightarrow\mathbb{R}^{2}$, $F_{h}:\mathbb{R}^{2}\times\mathcal{W}\rightarrow\mathbb{R}^{2}$. Here $\mathcal{U}=\left\{R_{u}\begin{bmatrix}\cos\left(\alpha)\right)\\ \sin\left(\alpha)\right)\end{bmatrix}\in\mathbb{R}^{2}:R_{u}\in\mathbb{R}_{+},\alpha\in[0,2\pi]\right\}$, and $\mathcal{W}=\left\{R_{w}\begin{bmatrix}\cos\left(\gamma)\right)\\ \sin\left(\gamma)\right)\end{bmatrix}\in\mathbb{R}^{2}:R_{w}\in\mathbb{R}_{+},\gamma\in[0,2\pi]\right\}$. Where $\mathbb{R}_{+}$ is the set of nonnegative real numbers. We let $w$ take different values within the set $\mathcal{W}$ with different probabilities. The models defined in (1) are known as constant position models [14].
@@
We assume that the robot and the moving obstacle are solid bodies with radius of $\frac{R}{2}$. This assumption simplifies the problem, but we can also handle the general case where the robot and obstacle have different sizes. Moreover, when the distance between the robot and the moving obstacle is less than or equal to $R$, i.e., $d\triangleq||r-h||_{2}\leq R$, the robot has collided with the moving obstacle. Additionally, when the distance between the robot and the target is less than or equal to $R$, i.e. $e\triangleq||r-t||_{2}\leq R$, the robot has reached the target.
@@ II-B Stochastic Infinite Horizon Optimal Control Problem
Assume that $r_{0}$ and $h_{0}$ are the initial positions of the robot and the moving obstacle, respectively. The problem is to find an optimal policy that enables the robot to reach the final state, i.e. the target, while trying to avoid colliding with the moving obstacle. The mathematical description of this problem is as follows:
@@
subject $\text{subject to}\quad\begin{cases}r_{k+1}=F_{r}\left(r_{k},u_{k}\right)\\ h_{k+1}=F_{h}\left(h_{k},w_{k}\right)\end{cases},\quad k\geq 0$
@@
where $\mathbb{E}[\cdot]$ denotes mathematical expectation, and where $f:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ is the so-called incremental cost. We define this function for a given target position $t$ as
$f(h,r)=\begin{cases}0,\quad||h-t||_{2}\leq R\\ \lambda(||h-t||_{2}-1)^{2}+\frac{1}{||h-r||_{2}+\epsilon},\quad||h-t||_{2}>R\end{cases}$ (3)
=$=\begin{cases}0,\quad e\leq R\\ \lambda(e-1)^{2}+\frac{1}{d+\epsilon},\quad e>R\end{cases}\triangleq\bar{f}(d,e)$
where $\bar{f}:\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$, and where $\lambda\in\mathbb{R}_{++}$ is a tuning parameter, and ϵ++italic-$\epsilon\in\mathbb{R}_{++}$ is a small number. Here, $\mathbb{R}_{++}$ is the set of positive real numbers. The motivation behind selecting this function is that when the robot reaches the target, i.e., $e\leq R$, the incurred cost is zero. However, when the robot gets close to the moving obstacle or is far from the target, the incurred cost is high. Note that when the robot reaches the target, then the incurred cost is zero, making the target position $t$ a so-called cost-free state. In other words, once the robot reaches the target, it will stay there without incurring any cost. This cost free state is well-known in the literature of stochastic shortest path problem [15],[16] and is also referred to as the destination [17].
@@ II-C Solution of Optimal Control
If we find a solution $V$ to this equation, the optimal policy is the minimizing argument in the right-hand side of the equation. Thus the optimal policy is a function of $(h,r)$. Note that, the Bellman equation in (4) must be solved for all $(h,r)$ and for a given target position $t$. This problem leads to the ’curse of dimensionality’ due to the large state space. We can use value iterations [11] to solve the Bellman equation in (4):
$V_{k+1}(h,r)=\min_{u\in\mathcal{U}}\mathbb{E}\left[f(h,r)+V_{k}\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$ (5)
where $V_{0}(h,r)=0,\forall h,r\in\mathbb{R}^{2}$. The limit of this sequence as $k$ goes to infinity satisfies the Bellman equation.
@@ III Geometric Symmetry
In this section we will show that the value function has a symmetry, and we will use this property to reduce the domain of the value function.
@@
Consider the two pairs, $(h_{1},r_{1})$ and $(h_{2},r_{2})$, in Fig. 1. If $d_{1}=d_{2}$, $e_{1}=e_{2}$, and $\theta_{1}=\pm\theta_{2}$, then these pairs are said to be symmetric around $t$
$x$$y$$\mathbf{t}$$\mathbf{r_{1}}$$\mathbf{r_{2}}$$\mathbf{h_{1}}$$\mathbf{h_{2}}$$\mathbf{e_{1}}$$\phi_{1}$$\mathbf{e_{2}}$$\phi_{2}$$\mathbf{d_{1}}$$\mathbf{d_{2}}$$\theta_{2}$$\theta_{1}$
@@
Consider two symmetric pairs $(h_{1},r_{1})$ and $(h_{2},r_{2})$. Let $u\in\mathcal{U}$, and $w\in\mathcal{W}$. Furthermore, assume that the probability density function of $w$ only depends on $R_{w}$, i.e. it is constant for each value of $\gamma$. In Fig. 2, two symmetric pairs $(h_{1},r_{1})$ and $(h_{2},r_{2})$, along with their possible future values, are depicted for the case where $R_{u}$ and $R_{w}$ have fixed values. Let $\mu(h,r,R_{u},\alpha)=\mathbb{E}\left[f(h,r)+V_{k}\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$, and $\eta(h,r,R_{u})=\min_{\alpha\in[0,2\pi]}\mu(h,r,R_{u},\alpha)$, where $\mu:\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}_{+}\times[0,2\pi]\rightarrow\mathbb{R}$, and $\eta:\mathbb{R}^{2}\times\mathbb{R}^{2}\times\mathbb{R}_{+}\rightarrow\mathbb{R}$. Then we can rewrite (5) as follows
$V_{k+1}(h,r)=\min_{R_{u}\in\mathbb{R}_{+}}\min_{\alpha\in[0,2\pi]}\mathbb{E}\left[f(h,r)+V_{k}\left(F_{h}(h,w),F_{r}(r,u)\right)\right]$ (6)
$=\min_{R_{u}\in\mathbb{R}_{+}}\min_{\alpha\in[0,2\pi]}\mu(h,r,R_{u},\alpha)=\min_{R_{u}\in\mathbb{R}_{+}}\eta(h,r,R_{u})$
For every fixed value of $R_{u}$, and a minimizing $\alpha_{1}$ corresponding to $r_{1}$, there is a minimizing $\alpha_{2}$ corresponding to $r_{2}$ such that $\mu(h_{1},r_{1},R_{u},\alpha_{1})=\mu(h_{2},r_{2},R_{u},\alpha_{2})$. In Fig. 2, the future values of these two points for a fixed value of $R_{u}$ are depicted by $r_{1}^{+}$ and $r_{2}^{+}$, respectively. Moreover, it holds that $\eta(h_{1},r_{1},R_{u})=\eta(h_{2},r_{2},R_{u})$. Then we optimize over $R_{u}$. In this way we have optimality over all $u\in\mathbb{R}^{2}$. Therefore, the minimum value of the right-hand side of (5) is the same for these two pairs, regardless of the values of $\phi_{1}$ and $\phi_{2}$. Hence for (5), we have
$V_{k+1}\left(h_{1},r_{1}\right)=$ (7)
$V_{k+1}\left(t+e_{1}\begin{bmatrix}\cos(\phi_{1})\\ \sin(\phi_{1})\end{bmatrix}+d_{1}\begin{bmatrix}\cos(\phi_{1}+\theta_{1})\\ \sin(\phi_{1}+\theta_{1})\end{bmatrix},t+e_{1}\begin{bmatrix}\cos(\phi_{1})\\ \sin(\phi_{1})\end{bmatrix}\right)=$
$V_{k+1}\left(t+e_{2}\begin{bmatrix}\cos(\phi_{2})\\ \sin(\phi_{2})\end{bmatrix}+d_{2}\begin{bmatrix}\cos(\phi_{2}+\theta_{2})\\ \sin(\phi_{2}+\theta_{2})\end{bmatrix},t+e_{2}\begin{bmatrix}\cos(\phi_{2})\\ \sin(\phi_{2})\end{bmatrix}\right)$
$=V_{k+1}\left(h_{2},r_{2}\right)$
where $e_{1}=e_{2}$, $d_{1}=d_{2}$, $\theta_{1}=\pm\theta_{2}$ and $\phi_{1},\phi_{2}$ are arbitrary. Hence $V_{k+1}$ can be considered to be a function of only $d$, $e$, and $\theta$. Now, we want to show that also the right-hand side of (5) can be expressed in terms of only $d$, $e$, and $\theta$.
$x$$y$$\mathbf{r}_{1}^{+}$$\mathbf{r}_{2}^{+}$$\mathbf{t}$$\mathbf{r_{1}}$$\mathbf{r_{2}}$$\mathbf{h_{1}}$$\mathbf{h_{2}}$$\mathbf{e}$$\phi_{1}$$\mathbf{e}$$\phi_{2}$$\mathbf{d}$$\mathbf{d}$$\theta_{2}$$\theta_{1}$
Figure 2: Geometrical Representation of Position of Robot, Obstacle, and Target for two symmetric pairs with all possible Future Values
Theorem 1
Let $(h,r)$ be a specific pair defined by $(d,e,\theta)$ as in Fig. 3, i.e.,
@@
$\begin{cases}e_{+}=F_{e}(e,u)\\ \theta_{+}=F_{\theta}(d,e,\theta,u,w)\\ d_{+}=F_{d}(d,\theta,u,w)\end{cases}$ (9)
where $F_{e}:\mathbb{R}_{+}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$, $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0,\pi]$, and $F_{d}:\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}$ are functions defined in the appendix.
Proof:
The proof is given in the appendix. ∎
$x$$y$$\mathbf{t}$$\mathbf{r}$$\mathbf{h}$$\mathbf{r}_{+}$$\mathbf{h}_{+}$$\mathbf{e}_{+}$$\mathbf{e}$$\mathbf{d}_{+}$$\mathbf{d}$$\theta$$\theta_{+}$
@@
We now define
$W_{k}(d,e,\theta)=V_{k}\left(t+e\begin{bmatrix}1\\ 0\end{bmatrix}+d\begin{bmatrix}\cos(\theta)\\ \sin(\theta)\end{bmatrix},t+e\begin{bmatrix}1\\ 0\end{bmatrix}\right)$ (10)
where $W_{k}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,2\pi]\rightarrow\mathbb{R}$. Then we can equivalently write (5) as
$W_{k+1}(d,e,\theta)$ (11)
$=\min_{u\in\mathcal{U}}\mathbb{E}\left[\bar{f}(d,e)+W_{k}\left(F_{d}(d,\theta,u,w),F_{e}(e,u),F_{\theta}(d,e,\theta,u,w)\right)\right]$
The limit of this recursion will define a solution to (4). This limit is denoted by $W(d,e,\theta)$, where $W:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,2\pi]\rightarrow\mathbb{R}$.
@@
In this section, we propose to solve the recursion for $W_{k}$ in (11) using fitted value iterations. To use fitted value iterations, we introduce vector-valued function $\varphi:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\rightarrow\mathbb{R}^{p}$, known as feature vector, and parameters $a_{k}\in\mathbb{R}^{p}$. Our goal is to approximate the function $W_{k}(d,e,\theta)$, with $\tilde{W}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathbb{R}^{p}\rightarrow\mathbb{R}$. One possibility is to use a linear regression model as follows:

### v2

@@ Abstract
Navigating a collision-free and optimal trajectory for a robot is a challenging task, particularly in environments with moving obstacles such as humans. We formulate this problem as a stochastic optimal control problem. Since solving the full problem is computationally demanding, we introduce a tractable approximation whose Bellman equation can be solved efficiently. The resulting value function is then incorporated as a terminal penalty in an online rollout framework. We construct a trade-off curve between safety and performance to identify an appropriate weighting between them, and compare the performance with other methods. Simulation results show that our method reduces the time to target by a factor of two compared to the control barrier function method, while maintaining a larger minimum distance to the obstacle.
@@ I INTRODUCTION
There are numerous applications of autonomous mobile robots that are discussed in the literature. In general, these applications are divided into two groups: indoor applications, such as delivering packages, cargo, and cleaning large buildings [1], and outdoor field robotics application [2]. Finding an optimal path for the robot to reach its destination is a crucial task in these applications. Path planning is one of the primary challenges that must be solved before mobile robots can autonomously navigate and explore complex environments [3].
Based on the nature of the environment that the robot is working in, the mobile path planning is divided into two classes, i.e., static environment and dynamic environment [1][5]. In the static environment, the obstacles are static and do not move. In the dynamic environment, the obstacles move during the planning task, while the robot is moving to reach its destination. The primary goal of path planning is to ensure safe, efficient, and collision-free navigation in both static and dynamic environments [6]. Path planning algorithms can be broadly categorized into several groups. The first group is classical algorithms, such as Dijkstra’s algorithm. The second group is heuristic-based approaches, e.g., the $A^{\star}$ algorithm. The third group is artificial intelligence-based approaches, such as reinforcement learning, fuzzy logic, and so on. The fourth group is hybrid-based approaches, which combine elements from all the above classes of path planning [1]. However, this is not the only way to classify path planning algorithms. In [5][7], and [8], other classifications for different approaches to path planning are also presented.
@@
In this paper, we consider a case in which there is a robot and a stochastically moving obstacle in the environment. Also, we assume that there is a static target, and we want to control the robot to reach this target. We formulate the problem of finding an optimal path as an infinite horizon stochastic optimal control problem. The associated Bellman equation can be solved using value iteration. Notably, for many problems, achieving an exact solution is simply infeasible due to what Bellman properly termed the ‘curse of dimensionality’ in [9, p. 94]. This problem refers to the exponential increase in the required computations as the problem size grows [10, p. 282]. In order to circumvent this problem, we exploit geometric symmetry and propose to use [11] to reduce the dimension of the domain of the value function. We then solve the resulting reduced Bellman equation using fitted value iteration. Fitted value iteration is a parametric approximation method in which we parameterize the value function as a regression model such that it closely satisfies the Bellman equation [12, p. 211].
To obtain a policy in real time, we employ the rollout approach. In this method, the control sequence over a finite horizon is optimized online while the cost-to-go of a base policy serves as a terminal penalty [13, 14], and [15, Chapter 4]. At each time step, only the first control input of the optimized sequence is applied, and the procedure is repeated at the next step, resulting in a receding horizon implementation that improves upon the base policy.
Path planning is inherently a multi-objective task [4]. In particular, two common and often competing objectives are safety and performance [17]. In this paper, we introduce two cost terms: one associated with safety and the other with performance, and seek an appropriate trade-off between these objectives.
We will compare our method with the Control Barrier Function (CBF) approach to stochastic discrete time systems [16], and a receding horizon $A^{\star}$ algorithm. Numerical results show that the proposed method can outperform the CBF, and the receding horizon $A^{*}$ methods and achieves a balanced trade-off between the competing objectives. More precisely, our method reduces the expected time to reach the target by approximately a factor of two compared to CBF, while ensuring a larger expected minimum distance from the obstacle.
@@
    We formulate an optimal path planning problem as a stochastic optimal control problem, that outperforms other shortest path formulations on average.
@@
    We use geometric symmetry to reduce the domain of the value function, enabling its representation with a reduced number of variables, which significantly lowers the computational cost of solving the Bellman equation.
@@ II Mathematical Model
Consider a problem where there is a robot, a stochastically moving obstacle, and a static target in $2$-dimensional space. The goal for the robot is to reach the target as fast as possible without colliding with the obstacle. In this section, we define the mathematical model of the dynamics and formulate an optimal control problem that will meet the objectives.
@@ II-A System Dynamics
Let $r_{k}=\begin{bmatrix}r_{k}^{x}&r_{k}^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ and $h_{k}=\begin{bmatrix}h_{k}^{x}&h_{k}^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ represent the positions of the robot and the moving obstacle at time $k$, respectively. Moreover, assume that $t=\begin{bmatrix}t^{x}&t^{y}\end{bmatrix}^{T}\in\mathbb{R}^{2}$ is the target position. The dynamics of the robot and the moving obstacle are described by the following equations:
@@
where $F_{r}:\mathbb{R}^{2}\times\mathcal{U}\rightarrow\mathbb{R}^{2}$, $F_{h}:\mathbb{R}^{2}\times\mathcal{W}\rightarrow\mathbb{R}^{2}$. The sets $\mathcal{U}$ and $\mathcal{W}$ represent all possible movements that the the robot, and the moving obstacle can take, respectively. Here $\mathcal{U}=\left\{R_{u}\begin{bmatrix}\cos\left(\alpha\right)&\sin\left(\alpha\right)\end{bmatrix}^{T}\in\mathbb{R}^{2}:R_{u}\in\{0,1\},\alpha\in\mathcal{A}_{1}\right\}$, and $\mathcal{W}=\left\{R_{w}\begin{bmatrix}\cos\left(\gamma\right)&\sin\left(\gamma\right)\end{bmatrix}^{T}\in\mathbb{R}^{2}:R_{w}\in\{0,1\},\gamma\in\mathcal{A}_{2}\right\}$, where $\mathcal{A}_{i}=\{\frac{q\pi}{n_{i}}\mid q=0,1,\cdots,2n_{i}-1\}$, and $n_{i},i=1,2$ are natural numbers. We can rewrite the set $\mathcal{W}$ as $\{w^{0},w^{1},\cdots,w^{2n_{2}}\}$, in which $w^{i}=\begin{bmatrix}\cos\left(\frac{i\pi}{n_{2}}\right)&\sin\left(\frac{i\pi}{n_{2}}\right)\end{bmatrix}^{T}$, for $0\leq i\leq 2n_{2}-1$, and $w^{i}=\begin{bmatrix}0&0\end{bmatrix}^{T}$ for $i=2n_{2}$. We let $w$ take different values within the set $\mathcal{W}$ with different probabilities. The models defined in (1) are known as constant position models [18].
@@
[…]
Assumption 2
We assume that $w$ is radially symmetric, i.e., its probability density function depends only on the magnitude $R_{w}$ and is constant for all orientations $\gamma$.
@@ II-B Stochastic Infinite Horizon Optimal Control Problem
Assume that $r_{0}$ and $h_{0}$ are initial positions of the robot and the moving obstacle, respectively. The problem is to find an optimal policy that enables the robot to reach the final state, i.e., the target, while trying to avoid colliding with the moving obstacle. The objectives are twofold: (i) minimize the expected time to reach the target, and (ii) maximize the minimum distance maintained between the robot and the moving obstacle. Obtaining the optimal solution to this problem is non-trivial. We formulate it as a stochastic optimal control problem with a specifically designed incremental cost. This incremental cost is expressed as a linear combination of two terms: the first term encourages the robot to reach the target as quickly as possible, while the second term penalizes situations in which the robot comes too close to the moving obstacle. The mathematical description of this problem is:
@@
[…]
where $\mathbb{E}[\cdot]$ denotes mathematical expectation, and where $f:\mathbb{R}^{2}\times\mathbb{R}^{2}\rightarrow\mathbb{R}_{+}$ is the so-called incremental cost. Here, $\mathbb{R}_{+}$ is the set of nonnegative real numbers. We define the function $f$ for a given target position $t$ as
[…]
where $\lambda\in[0,1]$ is a tuning parameter, and $\epsilon\in\mathbb{R}_{++}$ is a small number. Here, $\mathbb{R}_{++}$ is the set of positive real numbers. The motivation behind selecting this function is that when the robot reaches the target, i.e., $r-t||_{2}\leq R$, the incurred cost is zero. This makes the target position $t$ a so-called cost-free state. In other words, once the robot reaches the target, it will stay there without incurring any cost. This cost free state is well-known in the literature of Stochastic Shortest Path (SSP) problem [19, 20] and is also referred to as the destination [21]. Moreover, when the robot gets close to the moving obstacle or is far from the target, the incurred cost is high.
@@ II-C Solution of Optimal Control
If we find a solution $V$ to this equation, the optimal policy is the minimizing argument in the right-hand side of the equation. Thus the optimal policy is a function of $(h,r)$. Note that, the Bellman equation in (4) must be solved for all $(h,r)$ and for a given target position $t$. This problem leads to the ’curse of dimensionality’ due to the large state space.
@@ III Geometric Symmetry
In this section, we utilize the symmetry of the value function to reduce its domain and solve the Bellman equation in a lower dimensional state space.
@@
[…]
By defining these new variables, we can rewrite the incremental cost in (3) as
$f(h,r)=\begin{cases}0,\quad e\leq R\\ \lambda(e-R)^{2}+(1-\lambda)\frac{1}{d+\epsilon},\quad e>R\end{cases}\triangleq\bar{f}(d,e)$ (5)
where $\bar{f}:\mathbb{R}_{+}\times\mathbb{R}_{+}\rightarrow\mathbb{R}_{+}$.
Definition 2
Consider the two pairs, $(h_{1},r_{1})$ and $(h_{2},r_{2})$, in Fig. 1. If $d_{1}=d_{2}$, $e_{1}=e_{2}$, and $\theta_{1}=\pm\theta_{2}$, then these pairs are said to be symmetric around $t$ .
[…]
@@
The paper [11] introduces a framework that exploits symmetry to achieve dimensionality reduction in dynamic programming formulations of finite-horizon stochastic optimal control problems. In these settings, the inherent symmetry is shown to be preserved by both the optimal cost-to-go function and the corresponding optimal policy. It can be shown, using the result of [11], that the value function inherits this symmetry for the case of the stochastic shortest path problem under suitable assumptions. In other words, $V(h_{1},r_{1})=V(h_{2},r_{2})$ for all symmetric pairs $(h_{1},r_{1})$ and $(h_{2},r_{2})$. See Appendix A for the full proof.
This reduces the domain of the value function from $4$ variables $\begin{bmatrix}h&r\end{bmatrix}^{T}\in\mathbb{R}^{4}$ to $3$ scalar variables $d,e$, and $\theta$. See Appendix B for the full derivation of the reduced state formulation.
Using this symmetry property, we evaluate the Bellman equation only on a cross-section of symmetric states defined by
$\mathcal{C}=\left\{x=\begin{bmatrix}h&r\end{bmatrix}^{T}\in\mathbb{R}^{4}:r^{y}=t^{y},r^{x}>t^{x}\right\}$ (6)
Hence, the Bellman equation in reduced coordinates becomes:
$W(d,e,\theta)=\min_{u\in\mathcal{U}}\mathbb{E}\left[\bar{f}(d,e)+W\left(\bar{F}(d,e,\theta,u,w)\right)\right]$ (7)
where $\bar{F}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]$ is the system dynamics on $\mathcal{C}$. To define the function $\bar{F}$, let $(h,r)$ be a pair defined by $(d,e,\theta)$ on $\mathcal{C}$ as in Fig. 2, i.e.,
@@
[…]
where $F_{e}:\mathbb{R}_{+}\times\mathcal{U}\rightarrow\mathbb{R}_{+}$, $F_{\theta}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow[0,\pi]$, and $F_{d}:\mathbb{R}_{+}\times[0,\pi]\times\mathcal{U}\times\mathcal{W}\rightarrow\mathbb{R}_{+}$, and $\bar{F}=\begin{bmatrix}F_{d}&F_{\theta}&F_{d}\end{bmatrix}^{T}$, and where
See Appendix C for the full derivation of the reduced-state dynamics.
[…]
@@
We can use value iterations as in [12] to solve the Bellman equation in (7):
$W_{k+1}(d,e,\theta)=\min_{u\in\mathcal{U}}\mathbb{E}\left[\bar{f}(d,e)+W_{k}\left(\bar{F}(d,e,\theta,u,w)\right)\right]$ (11)
In the next section, we use fitted value iteration for this recursion.
@@
In fitted value iteration, we approximate the value function with a parameterized function. We select a set of sample states and adjust the parameters so that the Bellman equation is closely satisfied at those samples. The parameters are then updated iteratively, and this process repeats until they converge.
We introduce a vector-valued function $\varphi:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\rightarrow\mathbb{R}^{p}$, known as a feature vector, and parameters $a_{k}\in\mathbb{R}^{p}$. Our goal is to approximate the function $W_{k}(d,e,\theta)$, with $\tilde{W}:\mathbb{R}_{+}\times\mathbb{R}_{+}\times[0,\pi]\times\mathbb{R}^{p}\rightarrow\mathbb{R}$. One possibility is to use a linear regression model as follows: